WHen the numerator of a fraction is increased by 4 , the fraction increaces by 2/3 . What is the denominator of the fraction
Answer by venugopalramana(3286)

LET THE FRACTION BE A/B...N.R. IS INCREASED BY 4 SO IT IS A+4...ITS VALUE IS
(A+4)/B=A/B + 4/B......THIS IS 2/3 MORE THAN ORIGINAL FRACTION..SO
A/4 + 4/B = A/B + 2/3
SO 4/B = 2/3
2*B=4*3=12
B=12/2=6.......SO DENOMINATOR IS 6.